# Solving Football by Exploiting Equilibrium Structure of 2p0s Differential Games with One-Sided Information

### v3

@@ C.2 Preliminary lemmas
Let $\tilde{\phi}\in C^{2}$ be a test function touching $\tilde{V}$ at $(t,x,p)$, where $\tilde{V}$ is the non-revealing value. Lemma C.2 gives $|\tilde{\phi}(t+\tau,x+\tau f,p)-\tilde{V}(t+\tau,x+\tau f,p)|\leq C\tau^{2}$. Apply Lemma C.3 and use the fact that $\phi$ ($V$) is a convexification of $\tilde{\phi}$ ($\tilde{V}$) to get $|T_{\tau}[V]-T_{\tau}[\phi]|\leq C\tau^{2}$. ∎
Lemma C.5 (Consistency).
There exists a constant $C>0$ independent of $(t,x,p,\tau)$ such that
$|V(t,x,p)-T_{\tau}[V](t+\tau,x,p)|\leq C\tau^{2}$ (19)
for any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$.
Proof.
Let $\phi\in C^{2}$ be a test function touching $V$ at $(t_{0},x_{0},p_{0})$. Taylor expansion gives:
$T_{\tau}[\phi](t_{0}+\tau,x_{0},p_{0})=\phi(t_{0},x_{0},p_{0})+\tau\sum_{k=1}^{I}\lambda^{k}(\partial_{t}\tilde{\phi}(t_{0},x_{0},p^{k})+H^{k})+\epsilon_{\tau},$ (20)
where $\tilde{\phi}$ is the non-revealing value and $(u^{k},v^{k},\lambda^{k},p^{k})$ is the $k$th splitting point, so that $\phi(t_{0},x_{0},p_{0})=\sum_{k=1}^{I}\lambda^{k}\tilde{\phi}(t_{0},x_{0},p^{k})$. $H^{k}=\nabla_{x}\tilde{\phi}(t_{0},x_{0},p^{k})^{\top}f(x_{0},u^{k},v^{k})+\sum_{i=1}^{I}p^{k}[i]l_{i}(u^{k},v^{k})$. $\epsilon_{\tau}\leq C_{exp}\tau^{2}$.
First let $\phi$ and $\tilde{\phi}$ touch $V$ from above, i.e., $V\leq\phi\leq\tilde{\phi}$, $V(t_{0})=\phi(t_{0})$. From the viscosity property, we have $\partial_{t}\tilde{\phi}(t_{0},x_{0},p^{k})+H^{k}\geq 0$ for $k\in[I]$. Combining with Lem. C.4 to have
$T_{\tau}[V](t_{0}+\tau,x_{0},p_{0})\geq\phi(t_{0},x_{0},p_{0})-C\tau^{2}.$ (21)
Since $V(t_{0},x_{0},p_{0})=\phi(t_{0},x_{0},p_{0})$, we have
$V(t_{0},x_{0},p_{0})\leq T_{\tau}[V](t_{0}+\tau,x_{0},p_{0})+C\tau^{2}.$ (22)
We can similarly introduce $\phi^{\prime}$ and $\tilde{\phi}^{\prime}$ touching $V$ from below to have
$V(t_{0},x_{0},p_{0})\geq T_{\tau}[V](t_{0}+\tau,x_{0},p_{0})-C\tau^{2}.$ (23)
Lemma C.6 (Non-expansive Bellman backup).
Let $g$ and $h$ be two bounded functions defined on $0,T]\times\mathcal{X}\times\Delta(I)$. Then we have
$|T_{\tau}[g](t^{\prime},x,p)-T_{\tau}[h](t^{\prime},x,p)|\leq\sup_{x^{\prime}\in\mathcal{X},~p^{\prime}\in\Delta(I)}|g(t^{\prime},x^{\prime},p^{\prime})-h(t^{\prime},x^{\prime},p^{\prime})|$ (24)
for any $(t^{\prime},x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$.
Proof.
Let $\delta=\sup|g-h|$. Then $g\leq h+\delta$ and $h\leq g+\delta$. By monotonicity of $T_{\tau}$: $T_{\tau}[g]\leq T_{\tau}[h+\delta]$. It can be shown that $T_{\tau}[\cdot+\delta]=T_{\tau}[\cdot]+\delta$ for a constant $\delta$ since the value function shift adds directly. So, $T_{\tau}[g]\leq T_{\tau}[h]+\delta$. Similarly, $T_{\tau}[h]\leq T_{\tau}[g]+\delta$. Thus, $|T_{\tau}[g]-T_{\tau}[h]|\leq\delta$. ∎
Lemma C.7 (Value gap between $h$ and $h_{\tau}$).
Let $V(t,x,p)$ and $V_{\tau}(t,x,p)$ be the values of $h$ and $h_{\tau}$, respectively, at any $(t,x,p)\in[0,T]\times\mathcal{X}\times\Delta(I)$. We have
$|V(t,x,p)-V_{\tau}(t,x,p)|\leq C(T-t)\tau$
for some constant $C>0$ independent of $(t,x,p,\tau)$.
Proof.
Let $e(t,x,p)=V(t,x,p)-V_{\tau}(t,x,p)$ be the error function. Let $t_{k}=T-k\tau$ for $k\in[N]$, and $N=T/\tau$. Define $E_{k}=\sup_{x,p}|e(t_{k},x,p)|$. Note that $E_{0}=\sup_{x,p}|V(T,x,p)-V_{\tau}(T,x,p)|=0$ due to the terminal boundary.
Consider the error at time $t_{k}$ ($k\geq 1$):
$e(t_{k},x,p)$ $=V(t_{k},x,p)-V_{\tau}(t_{k},x,p)$
$=V(t_{k},x,p)-T_{\tau}[V_{\tau}](t_{k-1},x,p)\quad(\text{using }t_{k}+\tau=t_{k-1})$
Using the consistency bounds from Eq. 19:
$e(t_{k},x,p)$ $\leq(T_{\tau}[V](t_{k-1},x,p)+C\tau^{2})-T_{\tau}[V_{\tau}](t_{k-1},x,p)$
$=\left(T_{\tau}[V](t_{k-1},x,p)-T_{\tau}[V_{\tau}](t_{k-1},x,p)\right)+C\tau^{2}$
$e(t_{k},x,p)$ $\geq(T_{\tau}[V](t_{k-1},x,p)-C\tau^{2})-T_{\tau}[V_{\tau}](t_{k-1},x,p)$
$=\left(T_{\tau}[V](t_{k-1},x,p)-T_{\tau}[V_{\tau}](t_{k-1},x,p)\right)-C\tau^{2}$
From Lem. C.6, we have $|T_{\tau}[V](t_{k-1})-T_{\tau}[V_{\tau}](t_{k-1})|\leq E_{k-1}$. This implies:
$T_{\tau}[V](t_{k-1})-T_{\tau}[V_{\tau}](t_{k-1})$ $\leq E_{k-1}$
$T_{\tau}[V](t_{k-1})-T_{\tau}[V_{\tau}](t_{k-1})$ $\geq-E_{k-1}$
Substituting these into the bounds for $e(t_{k})$:
$e(t_{k},x,p)$ $\leq E_{k-1}+C\tau^{2}$
$e(t_{k},x,p)$ $\geq-E_{k-1}-C\tau^{2}$
Combining these gives:
$|e(t_{k},x,p)|\leq E_{k-1}+C\tau^{2}$
Since this holds for all $(x,p)$, we can take the supremum over $(x,p)$:
$E_{k}=\sup_{x,p}|e(t_{k},x,p)|\leq E_{k-1}+C\tau^{2}$
We have the recursion $E_{k}\leq E_{k-1}+C\tau^{2}$ with $E_{0}=0$, which leads to $E_{k}\leq kC\tau^{2}$. Use $k=(T-t_{k})/\tau$ to have
$E_{k}\leq\frac{T-t_{k}}{\tau}C\tau^{2}=C(T-t_{k})\tau$
Since $E_{k}=\|V(t_{k})-V_{\tau}(t_{k})\|_{\infty}$, we have shown that for any discrete time $t_{k}=T-k\tau$:
$\sup_{x,p}|V(t_{k},x,p)-V_{\tau}(t_{k},x,p)|\leq C(T-t_{k})\tau.$
@@
Our proof focuses on the primal game. The same technique can be applied to the dual game. First, it is easy to see $V(t,x,p)\leq\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$ because $\{\eta_{i,{\tau}}^{\dagger}\}$ is not necessarily NE in $G$. So we just need to prove
$\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)\leq V(t,x,p)+C(T-t)\tau,$ (25)
@@
To do so, we note that
$\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)\leq V_{\tau}(t,x,p).$ (26)
This is because in $G_{\tau}$, P2 moves after P1 and has an advantage. Then we just need to use Lem. C.7 to reach Eq. 25. ∎
@@ Appendix D Prediction Error of Value Approximation
Here we show that CAMS shares the same exponential error propagation as in standard approximate value iteration (AVI). The only difference from AVI is that the measurement error in CAMS comes from numerical approximation of the minimax problems rather than randomness in state transition and rewards. To start, let the true value be $V(t,x,p)$. Following Zanette et al. (2019), the prediction error $\epsilon^{bias}_{t}:=\max_{x,p}|\hat{V}_{t}(x,p)-V(t,x,p)|$ is affected by (1) the prediction error $\epsilon^{bias}_{t+\tau}$ propagated back from $t+\tau$, (2) the minimax error $\epsilon^{minmax}_{t}$ caused by limited iterations in solving the minimax problem at each collocation point: $\epsilon^{minmax}_{t}=\max_{(x,p)\in\mathcal{S}_{t}}|\tilde{V}(t,x,p)-V(t,x,p)|$, and (3) the approximation error due to the fact that $V(t,\cdot,\cdot)$ may not lie in the model hypothesis space $\mathcal{V}_{t}$ of $\hat{V}_{t}$: $\epsilon^{app}_{t}=\max_{x,p}\min_{\hat{V}_{t}\in\mathcal{V}_{t}}|\hat{V}_{t}(x,p)-V(t,x,p)|$.
@@
For simplicity, we will abuse the notation by using $x$ in place of $(x,p)$ and omit time dependence of variables when possible. In practice we consider $\hat{V}_{t}$ as neural networks that share the architecture and the hypothesis space. Note that $\hat{V}_{T}(\cdot)=V(T,\cdot)$ is analytically defined by the boundary condition and thus $\epsilon^{app}_{T}=\epsilon^{bias}_{T}=0$. To enable the analysis on neural networks, we adopt the assumption that $\hat{V}$ is infinitely wide and that the resultant neural tangent kernel (NTK) is positive definite. Therefore from NTK analysis (Jacot et al., 2018), $\hat{V}$ can be considered a kernel machine equipped with a kernel function $r(x^{(i)},x^{(j)}):=\langle\phi(x^{(i)}),\phi(x^{(j)})\rangle$ defined by a feature map $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d_{\phi}}$. Given training data $\mathcal{S}=\{(x^{(i)},V^{(i)})\}$, let $r(x)[i]:=r(x^{(i)},x)$, $R_{ij}:=r(x^{(i)},x^{(j)})$, $V_{\mathcal{S}}:=[V^{(1)},...,V^{(N)}]^{\top}$, $\Phi_{\mathcal{S}}:=[\phi(x^{(1)}),...,\phi(x^{(N)})]$, and $w_{\mathcal{S}}:=\Phi_{\mathcal{S}}(\Phi_{\mathcal{S}}^{\top}\Phi_{\mathcal{S}})^{-1}V_{\mathcal{S}}$ be model parameters learned from $\mathcal{S}$, then
@@ Appendix E Multigrid Algorithms and Results
Alg. 1 presents the $n$-level multigrid algorithm.
@@
In Fig. 6 we compare learned trajectories via the multigrid approach against the ground truth. The learned trajectories closely resemble the ground truth as P1 successfully concealing his payoff type until a critical time. In Fig. 6 we visualize the NE trajectories of P2 by solving the dual game.

### v4

@@ C.2 Preliminary lemmas
Let $\tau=T/L$, $\{\eta_{i,\tau}^{\dagger}\}$ be the random strategy associated with the feedback strategy $(u^{k},\alpha_{ki})$ in the game $G_{\tau}$, and $\zeta\in\mathcal{Z}$. Then, for any $k\in\{0,\dots,L\}$, $x\in\mathbb{R}^{d_{x}}$, and $p\in\Delta(I)$, claim that:
$\max_{\zeta\in\mathcal{Z}}J(t_{k},x,p;\{\eta_{i,\tau}^{\dagger}\},\zeta)\leq V_{\tau}(t_{k},x,p)+C(L-k)\tau^{2}$ (26)
We use backward induction to prove the equation (26). At $k=L$, the inequality holds trivially as there is no time left in the game and the error term goes to 0. Then, assume that (26) holds for $k+1$:
$\max_{\zeta\in\mathcal{Z}(t_{k+1})}J(t_{k+1},x,p;\{\eta_{i,\tau}^{\dagger}\},\zeta)\leq V_{\tau}(t_{k+1},x,p)+C(L-(k+1))\tau^{2}$ (27)
Now, to prove for any $k$, let us analyze what happens on $t_{k},t_{k+1})$. From the feedback given by ($\text{P}_{1}$), for a fixed $(x,p)$ we have:
$X_{T}^{t_{k},x,\{\eta_{i,\tau}^{\dagger}\},\zeta}=X_{T}^{t_{k+1},x^{j},\{\eta_{i,\tau}^{j\dagger}\},\zeta^{j}}\quad\text{with probability }\lambda^{j}p^{j}[i]/p[i],$
where
  • $x^{j}=X_{t_{k+1}}^{t_{k},x,u^{j},\zeta(u^{j})}$,
  • $\{\eta_{i,\tau}^{j\dagger}\}$ is the continuation of the random strategy from time $t_{k+1}$ associated with the same feedback but starting at $(t_{k+1},x^{j},p^{j})$,
  • $\zeta^{j}\in\mathcal{Z}(t_{k+1})$ is the continuation of the non-anticipative $\zeta$ after $u^{j}$ is chosen.
Then, using Lemma C.3, there is some $w^{j}\in\text{Co}f(x,u^{j},\mathcal{V})$ such that:
$|x^{j}-(x+\tau w^{j})|\leq C_{1}\tau^{2}.$ (28)
From the definition of $J$,
$J(t_{k},x,\{\eta_{i,\tau}^{\dagger}\},\zeta,p)$
$=\sum_{i=1}^{I}p[i]\mathbb{E}_{\{\eta_{i,\tau}^{\dagger}\}}\Big[g_{i}\Big(X_{T}^{t_{k},x,\{\eta_{i,\tau}^{\dagger}\},\zeta}\Big)+\int_{t_{k}}^{T}l_{i}\Big(X(s),\{\eta_{i,\tau}^{\dagger}\}(s),\zeta(s)\Big)\;ds\Big]$
$=\sum_{i=1}^{I}\sum_{j=1}^{I+1}p[i]\lambda^{j}p^{j}[i]/p[i]\mathbb{E}_{\{\eta_{i,\tau}^{j^{\dagger}}\}}\Big[g_{i}\Big(X_{T}^{t_{k+1},x,\{\eta_{i,\tau}^{j^{\dagger}}\},\zeta^{j}}\Big)+\int_{t_{k}}^{t_{k+1}}l_{i}(x(s),u^{j},\zeta(u^{j}))\;ds$
$\hskip 231.26378pt+\int_{t_{k+1}}^{T}l_{i}\Big(X(s),\{\eta_{i,\tau}^{j^{\dagger}}\}(s),\zeta^{j}(s)\Big)\;ds\Big]$
$\leq\sum_{j=1}^{I+1}\lambda^{j}\max_{\zeta^{\prime}\in\mathcal{Z}(t_{k+1})}\sum_{i=1}^{I}p^{j}[i]\mathbb{E}_{\{\eta_{i,\tau}^{j^{\dagger}}\}}\Big[g_{i}\Big(X_{T}^{t_{k+1},x,\{\eta_{i,\tau}^{j^{\dagger}}\},\zeta^{\prime}}\Big)+\int_{t_{k}}^{t_{k+1}}l_{i}(x(s),u^{j},\zeta(u^{j}))\;ds$
$\hskip 231.26378pt+\int_{t_{k+1}}^{T}l_{i}\Big(X(s),\{\eta_{i,\tau}^{j\dagger}\}(s),\zeta^{\prime}(s)\Big)\;ds\Big]$
Combining the above result with (27), we get:
$J(t_{k},x,\{\eta_{i,\tau}^{\dagger}\},\zeta,p)\leq\sum_{j=1}^{I+1}\lambda^{j}\Big[V_{\tau}(t_{k+1},x^{j},p^{j})+C(L-(k+1))\tau^{2}\Big].$
Finally, applying (28) and assuming $C_{2}$-Lipschitz continuity of $V_{\tau}$, we arrive at:
$J(t_{k},x,\{\eta_{i,\tau}^{\dagger}\},\zeta,p)\leq\sum_{j=1}^{I+1}\lambda^{j}\Big[V_{\tau}(t_{k+1},x+\tau w^{j},p^{j})+C_{2}|x^{j}-(x+\tau w^{j})|+$
$\hskip 289.07999ptC(L-(k+1))\tau^{2}\Big]$
$\hskip 36.135pt\leq\sum_{j=1}^{I+1}\lambda^{j}\Big[V_{\tau}(t_{k+1},x+\tau w^{j},p^{j})+C_{1}C_{2}\tau^{2}+C(L-(k+1))\tau^{2}\Big]$
$\hskip 36.135pt\leq\sum_{j=1}^{I+1}\lambda^{j}\max_{w\in\text{Co}f(x,u^{j},\mathcal{V})}\Big[V_{\tau}(t_{k+1},x+\tau w,p^{j})+C_{1}C_{2}\tau^{2}+C(L-(k+1))\tau^{2}\Big]$
$\hskip 36.135pt\leq V_{\tau}(t_{k},x,p)+(C_{1}C_{2}+2)\tau^{2}+C(L-(k+1))\tau^{2}.$
The term $2\tau^{2}$ arises from two approximation steps: first $\tau^{2}$ from selecting an approximately optimal convex decomposition in the definition of $V_{\tau}(t_{k},x,p)$, and the second $\tau^{2}$ from choosing $u^{j}$ to be only $\tau^{2}$-optimal for $\min_{u}\max_{v}(\cdot)$ so that they are Borel measurable with respect to $(x,p)$.
Then, choosing $C=(C_{1}C_{2}+2)$ satisfies (26).
Since we have, $L=T/\tau$, and $t_{k}=k\tau$. We get
$\max_{\zeta\in\mathcal{Z}}J(t_{k},x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)\leq V_{\tau}(t_{k},x,p)+C(T-t_{k})\tau$
Note that compared with the proof provided in cardaliaguet2009numerical Theorem 4.1, our proof incorporates instantaneous costs and clarified the stage-wise $\mathcal{O}(\tau)$ error in (27).
@@
Our proof focuses on the primal game. The same technique can be applied to the dual game. First, it is easy to see $V(t,x,p)\leq\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$ because $\{\eta_{i,{\tau}}^{\dagger}\}$ is not necessarily NE in $G$.
Next, we need to show
$0\leq\max_{\zeta\in\mathcal{Z}}J(t,x,p;\{\eta_{i,{\tau}}\}^{\dagger},\zeta)-V(t,x,p)\leq\epsilon$ (29)
@@
Let us fix some $t^{\prime}$, $x^{\prime}$, and $p^{\prime}$, and some $k$ such that $t^{\prime}\in[t_{k-1},t_{k})$. Then,
$\max_{\zeta\in\mathcal{Z}_{r}(t^{\prime})}J(t^{\prime},x^{\prime},p^{\prime};\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$ $=\max_{\zeta\in\mathcal{Z}(t^{\prime})}J(t^{\prime},x^{\prime},p^{\prime};\{\eta_{i,{\tau}}^{\dagger}\},\zeta)$
[…]
where $\{\eta_{i,\tau}^{x\dagger}\}$ is the strategy at $(t_{k},x,p^{\prime})$, $F$ is a bound on $f$, and $l||_{\infty}$ term is due to bounded instantaneous cost. Hence, from Lemma C.4
$\max_{\zeta\in\mathcal{(}t^{\prime})Z}J(t^{\prime},x^{\prime},p^{\prime},\{\eta_{i,\tau}\}^{\dagger},\zeta)\leq\max_{|x-x^{\prime}|\leq F\tau}V_{\tau}(t_{k},x,p^{\prime})+\mathcal{O}(\tau)$
[…]
Therefore, for any $\epsilon>0$ there exists $\tau^{*}>0$ such that for all $\tau\in(0,\tau^{*}]$ inequality (29) holds. ∎
@@ Appendix D Prediction Error of Value Approximation
Here we show that CAMS shares the same exponential error propagation as in standard approximate value iteration (AVI). The only difference from AVI is that the measurement error in CAMS comes from numerical approximation of the minimax problems rather than randomness in state transition and rewards. To start, let the true value be $V(t,x,p)$. Following zanette2019limiting, the prediction error $\epsilon^{bias}_{t}:=\max_{x,p}|\hat{V}_{t}(x,p)-V(t,x,p)|$ is affected by (1) the prediction error $\epsilon^{bias}_{t+\tau}$ propagated back from $t+\tau$, (2) the minimax error $\epsilon^{minmax}_{t}$ caused by limited iterations in solving the minimax problem at each collocation point: $\epsilon^{minmax}_{t}=\max_{(x,p)\in\mathcal{S}_{t}}|\tilde{V}(t,x,p)-V(t,x,p)|$, and (3) the approximation error due to the fact that $V(t,\cdot,\cdot)$ may not lie in the model hypothesis space $\mathcal{V}_{t}$ of $\hat{V}_{t}$: $\epsilon^{app}_{t}=\max_{x,p}\min_{\hat{V}_{t}\in\mathcal{V}_{t}}|\hat{V}_{t}(x,p)-V(t,x,p)|$.
@@
For simplicity, we will abuse the notation by using $x$ in place of $(x,p)$ and omit time dependence of variables when possible. In practice we consider $\hat{V}_{t}$ as neural networks that share the architecture and the hypothesis space. Note that $\hat{V}_{T}(\cdot)=V(T,\cdot)$ is analytically defined by the boundary condition and thus $\epsilon^{app}_{T}=\epsilon^{bias}_{T}=0$. To enable the analysis on neural networks, we adopt the assumption that $\hat{V}$ is infinitely wide and that the resultant neural tangent kernel (NTK) is positive definite. Therefore from NTK analysis (jacot2018neural), $\hat{V}$ can be considered a kernel machine equipped with a kernel function $r(x^{(i)},x^{(j)}):=\langle\phi(x^{(i)}),\phi(x^{(j)})\rangle$ defined by a feature map $\phi:\mathcal{X}\rightarrow\mathbb{R}^{d_{\phi}}$. Given training data $\mathcal{S}=\{(x^{(i)},V^{(i)})\}$, let $r(x)[i]:=r(x^{(i)},x)$, $R_{ij}:=r(x^{(i)},x^{(j)})$, $V_{\mathcal{S}}:=[V^{(1)},...,V^{(N)}]^{\top}$, $\Phi_{\mathcal{S}}:=[\phi(x^{(1)}),...,\phi(x^{(N)})]$, and $w_{\mathcal{S}}:=\Phi_{\mathcal{S}}(\Phi_{\mathcal{S}}^{\top}\Phi_{\mathcal{S}})^{-1}V_{\mathcal{S}}$ be model parameters learned from $\mathcal{S}$, then
@@ Appendix E Multigrid Algorithms and Results
Fig. 7 shows a typical 2-level multigrid (Full-Approximation Scheme or FAS) approach. As discussed, FAS has four steps, namely: (1) restriction of the fine-grid approximation and its residual into the coarse grid (red arrows in Fig. 7); (2) computation of the coarse-grid solution by incorporating restricted fine-grid residuals; (3) computation of the coarse-grid correction; and finally, (4) prolongation of the coarse-grid correction to the fine-grid (shown by the blue arrows in Fig. 7). This can be further extended to $n$-level multigrid by recursively reducing the coarse-grid size until the desired coarsest grid is reached. Alg. 1 presents the $n$-level multigrid algorithm.
Figure 7: Illustration of 2-level multigrid method.
@@
In Fig. 8 we compare learned trajectories via the multigrid approach against the ground truth. The learned trajectories closely resemble the ground truth as P1 successfully concealing its payoff type until a critical time. In Fig. 8 we visualize the NE trajectories of P2 by solving the dual game.